OpenStudy (anonymous):

Find an expression for a cubic function f if f(1)=24 and f(-3)=f(0)=f(2)=0

OpenStudy (anonymous):

Refer to the Mathematica 9 plot and solution attached.

OpenStudy (kirbykirby):

You can do this without Mathematica...unless your question required software. A cubic function can be generally expressed as \(\large f(x) =ax^3+bx^2+cx+d\) With the values you have, you can essentially create a system of 4 equations and 4 unknowns: \[\large f(1)=a(1)^3+b(1)^2+c(1)+d=24\\ \large f(-3)=a(-3)^3+b(-3)^2+c(-3)+d=0 \\ \large f(0)=a(0)^3+b(0)^2+c(0)+d=0 \\ \large f(2)=a(2)^3+b(2)^2+c(2)+d=0 \] Notice how the 3rd question though immediately tells you that \(d=0\), effectively giving you now a system of 3 equations to solve for a, b, c

OpenStudy (kirbykirby):

3rd equation*